If Y is the midpoint of WZ, YZ = x + 3 and WZ = 3x-4 - determine the length of WZ
Answer by stanbon(75887)

If Y is the midpoint of WZ, YZ = x + 3 and WZ = 3x-4 -
determine the length of WZ
------
Draw the picture::
W..........Y..........Z
------
WY = YZ
WY = x+3
-----
WZ = 2(x+3) = 2x+6
-----
3x-4 = 2x+6
x = 10
------
WZ = 2*10+6 = 26
---------
Cheers,
Stan H.
-------------